UT 20150509 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson (all remote)
Observer AO: DMiller (LBTO, remote)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Continue AO-com depending on observing conditions (most likely not so good)

Summary

- Lost most of the night due to high humidity

- Lost some time due to TCS problems

- Some (very few) ghost data taken

Details

All times are UT
Files are luci2.20150509.NNNN.fits

02:10 Sunset, wind close to the limits wait a bit, clear, T = -2deg, Hum = 70% (rising)

02:35 Wind now 25mins below limit, Steve opens, points and collimates

Seeing is poor 1.5-3" due to low-pressure system pumping cold air above us

03:05 Problems with shell

04:00 doing Ghosts at K=9mag star, 30000cts, FWHM ~0.12", seeing at guider about 1.2"
Correction: 800Hz, Bin = 2, 100modes

04:26 Image quality at medium observing conditions, NGC 4147

Failed, TCS died, shell ripped (IT 5559)

seeing too bad for NGC 4147, couldn't collimate on a 14.9mag star

04:45 Test PSF-stability long/short term at intermediate conditions on 3C 294
AO-ref star has R=11.3, Bin=2, preset succesfull, run-AO failed

05:00 Humidity above 90% now, wind is rising again

05:10 Redo 3C 294, 1.4mag extinction => 4mag extinction, run-AO failed
redo it again, nothing else really useful to be done given the deteriorate observing conditions

TCS is having problem with its internal communication again

05:35 Humidity above 95%, closing

10:30 Hum = 100%, T= -7deg, high wind, call it a night

-- %USERSIG{JochenHeidt - 2015-05-09}%
Topic revision: r3 - 09 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback